The voltage applied to a circuit is v=100 sin(wt+30 º) and current flowing in the circuit is i= 15 sin(wt+60 º). induced emf. Determine: 1. Impedance 2.Phase angle between voltage and current 3.Active Power 4.Power Factor.

1
Expert's answer
2021-06-08T09:35:07-0400

Gives

"V=100sin(wt+30\u00b0)"

"I=15sin(wt+60\u00b0)"

Impedance (Z)

"Z\sqrt{R^2+{(X_L-X_c})^2}"

"X_L=X_c=0"

Then

Z=R="\frac{100}{15}=6.66\Omega"

Phase difference


"∆\phi=(wt+60-wt+30)=30°"

Active power

"P=\frac{IV}{2}cos\phi"

Put value

"P=\frac{100\times15}{2}cos30°"

P=650 W

Power factor

"Cos\phi =\frac{R}{Z}"

"R=\frac{V}{I}"

"R=\frac{100}{15}"

"Cos\phi=\frac{6.66}{6.66}=1"

"\phi=cos^{-1}1=0°"

Power factor"\phi=0°"